# MNSC-0001-2024.00

### v1

@@ 4.1 Mixing Time of the Stochastic Process
4.2 Impact on Global Average Risk
In Section 3.3, we established that under Section 3.3, the marginal failure probabilities of individual nodes increase over time, as shown in Section 3.3. We then used this to establish the monotonicity of $\mathbb{E}\mathopen{}\mathclose{{\left[\mathcal{L}(\mathbf{X}^{t})}}\right]$. However, it is not immediately clear whether this monotonicity extends to the full distribution of the global average risk defined over the stochastic process $(\mathbf{X}^{t})_{t\in\mathbb{N}}$. The rest of this section focuses on establishing that the distribution of global average risk is monotone (i.e. stochastically dominated) with respect to $t\in\mathbb{N}$, establishing the impact of network risk propagation to long-term financial insolvency.
Our first result formalizes the intuition that, under Section 3.3, the probability of the global failure loss exceeding a given threshold is non-decreasing over time. The proof relies on constructing a coupling argument that establishes a form of stochastic dominance across time steps. This result confirms that as failures accumulate, they are unlikely to be reversed by the system dynamics alone. This reinforces the insight that traditional models, which often assume failures are independent, may underestimate long-term risk exposure when network dependencies are present.
{restatable}
theoremGARIncrease Under Assumption 3.3, we have that for any $\varepsilon\in\mathbb{R}_{\geq 0}$ and for all $t\in\mathbb{N}$
$\mathbb{P}\mathopen{}\mathclose{{\left(\mathcal{L}(\mathbf{X}^{t+1})>\varepsilon}}\right)\geq\mathbb{P}\mathopen{}\mathclose{{\left(\mathcal{L}(\mathbf{X}^{t})>\varepsilon}}\right).$ (14)
As a result,
$\mathbb{P}\mathopen{}\mathclose{{\left(\mathcal{L}(\mathbf{X}^{\infty})>\varepsilon}}\right)\geq\mathbb{P}\mathopen{}\mathclose{{\left(\mathcal{L}(\mathbf{X}^{0})>\varepsilon}}\right).$ (15)
Before presenting the proof we note that it leverages several results from muller2002comparison, which we include in Section 11 for completeness.
Proof 4.5
To prove Eq. 14, we first show that for any $i\in\mathcal{V}$ and $t\in\mathbb{N}$, $X_{i}^{t}\preceq_{\text{st}}X_{i}^{t+1}$ by using Strassen’s Theorem (Roch_2024, Theorem 4.2.11). Then, we verify that the vectors $\mathbf{X}^{t}$ and $\mathbf{X}^{t+1}$ are coupled using the same source of randomness, and thus share a common copula. Therefore, we can conclude by Theorem 11.1 (appendix) that $\mathbf{X}^{t}\preceq_{\text{st}}\mathbf{X}^{t+1}$ for all $t\geq 0$. Finally, we use the fact that $\mathcal{L}(\cdot)$ is a monotone increasing function and apply Theorem 11.2 (appendix) to conclude that $\mathcal{L}(\mathbf{X}^{t})\preceq_{\text{st}}\mathcal{L}(\mathbf{X}^{t+1})$, which is equivalent to Eq. 14.
Thus, we only need to show that $X_{i}^{t}\preceq_{\text{st}}X_{i}^{t+1}$ and that the coupling inducing the ordering is defined through a common copula. First, note that $\{0,1\}^{n}$ with the component-wise ordering is a finite poset. By Strassen’s Theorem, $X_{i}^{t}\preceq_{\text{st}}X_{i}^{t+1}$ if and only if there exists a coupling $(\tilde{X}_{i}^{t},\tilde{X}_{i}^{t+1})$ of $(X_{i}^{t},X_{i}^{t+1})$ such that $\tilde{X}_{i}^{t}\leq\tilde{X}_{i}^{t+1}$ almost surely under the coupling.
We now turn our efforts to constructing such a coupling. Recall $m_{i}^{t}\triangleq\mathbb{E}[X_{i}^{t}]$, and define for each node $i\in\mathcal{V}$, $U_{i}\overset{iid}{\sim}\text{Unif}(0,1)$. For each $t\geq 0$, we define the coupled process:
$\tilde{X}_{i}^{t}=\mathbbm{1}\{U_{i}\leq m_{i}^{t}\},$
which is a valid coupling since $\tilde{X}_{i}^{t}\sim\text{Ber}(m_{i}^{t})$, matching the marginal distribution of $X_{i}^{t}$. By Section 3.3, we know $m_{i}^{t}\leq m_{i}^{t+1}$. Thus, when $U_{i}\leq m_{i}^{t}\leq m_{i}^{t+1}$, we get $\tilde{X}_{i}^{t}=1=\tilde{X}_{i}^{t+1}$. Similarly, when $U_{i}>m_{i}^{t}$, then $\tilde{X}_{i}^{t}=0$, and $\tilde{X}_{i}^{t+1}\in\{0,1\}$. Therefore, for every $i$ and $t\geq 0$, we have
$\tilde{X}_{i}^{t}\leq\tilde{X}_{i}^{t+1}\quad\text{a.s. under the coupling.}$
By Strassen’s Theorem, this implies
$X_{i}^{t}\preceq_{\text{st}}X_{i}^{t+1}.$
Moreover, because this coupling uses the same uniform random vector $\mathbf{U}=(U_{1},\dots,U_{n})$ across all coordinates $i$, the vectors $\mathbf{X}^{t}$ and $\mathbf{X}^{t+1}$ are constructed using the same dependence structure. Specifically, since the $U_{i}$ are independent, this induces the product copula $C(u_{1},\dots,u_{n})=\prod_{i=1}^{n}u_{i}$ for both $\mathbf{X}^{t}$ and $\mathbf{X}^{t+1}$. Therefore, we conclude by Theorem 11.1 that $\mathbf{X}^{t}\preceq_{\text{st}}\mathbf{X}^{t+1}.$
Finally, since $\mathcal{L}(\mathbf{X}^{t})=\sum_{i=1}^{n}\beta_{i}X_{i}^{t}$ is a component-wise increasing function and $\beta_{i}>0$ for all $i$, Theorem 11.2 ensures that
$\mathcal{L}(\mathbf{X}^{t})\preceq_{\text{st}}\mathcal{L}(\mathbf{X}^{t+1}).$
That is, for every $\varepsilon\in\mathbb{R}$ and for all $t\geq 0$,
$\mathbb{P}\mathopen{}\mathclose{{\left(\mathcal{L}(\mathbf{X}^{t+1})>\varepsilon}}\right)\geq\mathbb{P}\mathopen{}\mathclose{{\left(\mathcal{L}(\mathbf{X}^{t})>\varepsilon}}\right),$
which shows Eq. 14. Eq. 15 follows since for each $\varepsilon\in\mathbb{R}$ the sequence $\{\mathbb{P}(\mathcal{L}(\mathbf{X}^{t})>\varepsilon)\}_{t\in\mathbb{N}}$ is non-decreasing. Moreover, by Fig. 4, we have $\mathcal{L}(\mathbf{X}^{t})\to\mathcal{L}(\mathbf{X}^{\infty})$ almost surely, which implies convergence in distribution. Hence, for every continuity point $\varepsilon$ of the limiting distribution, we obtain
$\lim_{t\to\infty}\mathbb{P}\bigl(\mathcal{L}(\mathbf{X}^{t})>\varepsilon\bigr)=\mathbb{P}\bigl(\mathcal{L}(\mathbf{X}^{\infty})>\varepsilon\bigr),$
and in particular, $\mathbb{P}\bigl(\mathcal{L}(\mathbf{X}^{0})>\varepsilon\bigr)\leq\mathbb{P}\bigl(\mathcal{L}(\mathbf{X}^{\infty})>\varepsilon\bigr).$
Theorem 4.2 shows that under Assumption 3.3, the entire distribution of global financial loss stochastically dominates the independent failure model. That is, not only does the expected loss increase over time (as in Corollary 8.10), but the quantiles also become more severe. This has strong implications for surety organizations, as they must retain financial reserves to absorb losses during rare but extreme events. If such tail risks are systematically underestimated (as in $\mathcal{L}(\mathbf{X}^{0})$ by ignoring network spillovers), then reserve requirements will be miscalibrated.
@@ 5.1 Description of Surety Network
See Fig. 6 for a visualization of the adjacency matrix of our surety network. We emphasize that the graph is mostly bipartite, with the exception of a small number of intermediaries. We further note that the connectivity pattern in the graph is sparse: among the 35,483 nodes in the giant connected component the edge density is around 0.009%, representing 56,707 contracts. The nature of the graph also reveals interesting structural properties — it is acyclic with a directed diameter of seven. This acyclic structure has important implications for our modeling: our stochastic failure-propagation process converges to its stationary distribution in at most seven time steps (see Definition 4.4). Lastly, we note some degree of heterogeneity in degree distribution across nodes, witnessed in Fig. 7.
@@ 5.2 Case Study
Turning to the upstream neighbors’ risk distribution (Fig. 9, right), we observe substantial heterogeneity: some neighbors have risk scores lower than node 39466’s idiosyncratic value $r_{39466}=0.00182$, while others are considerably higher. While some neighbors are less risky, when incorporating these exposures, the induced limiting failure probability increases to $m_{39466}=0.00264$, corresponding to a relative increase of roughly  $45\%$. This amplification illustrates how even a moderately risky intermediary, once identified through the $u_{39466}$ centrality metric, can see its effective default probability substantially elevated due to network position and contracting relationships.
@@ 5.3 Impact of network effects on global financial loss
Next we begin to illustrate the impact of network risk on the global average financial loss in the surety network. We start by recalling in Sections 3.3 and 4.2 that under Section 3.3, both the average risk and the tail probability mass of the loss distribution increases. However, this assumption is not always satisfied in our surety network (see more discussion on this in Section 10.3).
@@
Despite this, we observe that for a nontrivial subset of nodes $i\in\mathcal{V}$, the expected marginal loss probability $m_{i}$ exceeds their idiosyncratic risk scores $r_{i}$. This effect is visible in Fig. 10, which compares $r_{i}$ (x-axis) and $m_{i}$ (y-axis), with numerous points above the $y=x$ line. Moreover, as a result of this, the expected aggregate loss $\mathbb{E}\mathopen{}\mathclose{{\left[\mathcal{L}(\mathbf{X}^{\infty})}}\right]=\beta^{\top}\mathbf{m}$ increases from the expected aggregate loss under the independent failure model $\mathbb{E}\mathopen{}\mathclose{{\left[\mathcal{L}(\mathbf{X}^{0})}}\right]=\beta^{\top}\mathbf{r}$ by 1.89%. This emphasizes that failing to account for network interference causes downstream risk to be under-estimated. While 1.89% might seem like a mild value, we emphasize that the units for these are on the order of hundreds of millions of dollars, so this increase is roughly on the order of 2 million dollars of underestimated risk. See Fig. 10 where we also include a box plot of $\beta_{i}m_{i}$ vs $\beta_{i}r_{i}$ across nodes $i\in\mathcal{V}$. Here we observe not only does the average increase, but $\beta_{i}m_{i}$ is more right-skewed. This again emphasizes the ability for network risk to destabilize the network.
@@ 5.4 Robustness to choice of $\alpha_{i}$
We close out by testing the robustness of our earlier empirical insights with respect to the choice of $\alpha$, where the baseline specification set $\alpha_{i}=0.25$ for all intermediaries $i$ in the network. In Fig. 11, we report the quantiles of $\mathcal{L}(\mathbf{X}^{\infty})$ as a function of $\alpha\in[0,1]$, applied uniformly across all intermediaries. Across all specifications, the outcomes remain monotone in $\alpha$: greater exposure consistently increases systemic risk. This finding is noteworthy because the formal monotonicity condition (Section 3.3) does not hold in our empirical network. If it did, Section 4.2 would guarantee this result; the fact that monotonicity emerges regardless highlights the robustness of our conclusions.

### v2

@@ 5.1 Description of Surety Network
See Fig. 6 for a visualization of the adjacency matrix of our surety network. We emphasize that the graph is mostly bipartite, with the exception of a small number of intermediaries. We further note that the connectivity pattern in the graph is sparse: among the 35,483 nodes in the giant connected component the edge density is around 0.009%, representing 56,707 contracts. The nature of the graph also reveals interesting structural properties, it is acyclic with a directed diameter of seven. This acyclic structure has important implications for our modeling since our stochastic failure-propagation process converges to its stationary distribution in at most seven time steps (see Definition 4.4). Lastly, we note some degree of heterogeneity in degree distribution across nodes, witnessed in Fig. 7.
@@ 5.2 Case Study
Turning to the upstream neighbors’ risk distribution (Fig. 9, right), we observe substantial heterogeneity. Some neighbors have risk scores lower than node 39466’s idiosyncratic value $r_{39466}=0.00182$, while others are considerably higher. While some neighbors are less risky, when incorporating these exposures, the induced limiting failure probability increases to $m_{39466}=0.00264$, corresponding to a relative increase of roughly  $45\%$. This amplification illustrates how even a moderately risky intermediary, once identified through the $u_{39466}$ centrality metric, can see its effective default probability substantially elevated due to network position and contracting relationships.
@@ 5.3 Impact of network effects on global financial loss
Next we begin to illustrate the impact of network risk on the global average financial loss in the surety network. We start by recalling in Section 3.3 that under Section 3.3 the average risk increases. However, this assumption is not always satisfied in our surety network (see more discussion on this in Section 10.3).
@@
Despite this, we observe that for a nontrivial subset of nodes $i\in\mathcal{V}$, the expected marginal loss probability $m_{i}$ exceeds their idiosyncratic risk scores $r_{i}$. This effect is visible in Fig. 10 (left), which compares $r_{i}$ (x-axis) and $m_{i}$ (y-axis), with numerous points above the $y=x$ line. Moreover, as a result of this, the expected aggregate loss $\mathbb{E}\mathopen{}\mathclose{{\left[\mathcal{L}(\mathbf{X}^{\infty})}}\right]=\beta^{\top}\mathbf{m}$ increases from the expected aggregate loss under the independent failure model $\mathbb{E}\mathopen{}\mathclose{{\left[\mathcal{L}(\mathbf{X}^{0})}}\right]=\beta^{\top}\mathbf{r}$ by 1.89%. This emphasizes that failing to account for network interference causes downstream risk to be under-estimated. While 1.89% might seem like a mild value, we emphasize that the units for these are on the order of hundreds of millions of dollars, so this increase is roughly on the order of 2 million dollars of underestimated risk. See Fig. 10 (right) where we also include a box plot of $\beta_{i}m_{i}$ vs $\beta_{i}r_{i}$ across nodes $i\in\mathcal{V}$. Here we observe not only does the average increase, but $\beta_{i}m_{i}$ is more right-skewed. This again emphasizes the ability for network risk to destabilize the network.
@@ 5.4 Robustness to choice of $\alpha_{i}$
We close out by testing the robustness of our earlier empirical insights with respect to the choice of $\alpha$, where the baseline specification set $\alpha_{i}=0.25$ for all intermediaries $i$ in the network. In Fig. 11, we report the quantiles of $\mathcal{L}(\mathbf{X}^{\infty})$ as a function of $\alpha\in[0,1]$, applied uniformly across all intermediaries. Across all specifications, the outcomes remain monotone in $\alpha$ as greater exposure consistently increases systemic risk. This finding is noteworthy because the formal monotonicity condition (Section 3.3) does not hold in our empirical network.